OpenStudy (anonymous):

Let a be any positive number then: (-a)^(1/2) =(-a)^0.5 =(-a)^(0.48 + 0.02) =(-a)^(0.48) * (-a)^(0.02) =(-a)^(48/100) * (-a)^(2/100) = [{(-a)^(48)}^(1/100)] * [{(-a)^(2)}^(1/100}] = [{(a)^(48)}^(1/100)] * [{(a)^(2)}^(1/100}] =(a)^(0.48) * (a)^(0.02) =a^0.5 =a^(1/2) [WHAT IS WRONG]

OpenStudy (anonymous):

(-a)^(1/2) Using these can you tell me which ones are right?

OpenStudy (experimentx):

\[ -a = \sqrt{(-a)^2} = \sqrt{a^2} = a \] tell me what's wrong here? I'll tell you what's wrong there.

OpenStudy (anonymous):

I dont think I have squared and undid it again. I have just divided the number into different parts.

OpenStudy (experimentx):

(-a)^(48) = (a)^(48) <--- this is what you did

OpenStudy (anonymous):

Yes that is what I did. Isn't (-a)^(48) = (a)^(48) true

OpenStudy (experimentx):

you ignored 1/100 negative number do not love fractional power

OpenStudy (experimentx):

it seems that i might have mistaken ... i'll update it a bit later.